OpenStudy (anonymous):

the average of 15 numbers is 7. Angelo is adding these numbers and mistakenly reads 6 of the numbers as 9 instead of 7,what average will he get??..help plz..

OpenStudy (harsimran_hs4):

total numbers are 15 so every time he takes 9 instead of 6 he adds 3 extra to the previous total so previous total = \[15 \times 7\] and extra is 3 * 7 so find new total and divide by 15 any problem persists please ask

OpenStudy (harsimran_hs4):

*sorry he reads 7 instead of 9 so he adds 2

OpenStudy (anonymous):

sum/15=7 sum=15*7=105 "mistakenly reads 6 of the numbers as 9 instead of 7" => 6*2=12 Add 12 to original sum and calculate the average.